# Uncertainty in predicting the stochastic gravitational wave background from compact binary coalescences

### v2

@@ III.2 BBH spin models
We investigate the impact of different spin models on the BBH background energy spectrum and compare them to the non-spinning fiducial models. The GWTC-3 population analysis considered two spin models [11]. The first, referred to as the default model, treats spin magnitudes and orientations independently. Spin magnitudes for both black holes are drawn from a Beta distribution, while tilt angles, defined as the angle between each black hole’s spin and the orbital angular momentum, are drawn from a mixture of an isotropic distribution and a preferentially aligned component, modeled by a truncated Gaussian distribution [79]. The second, known as the Gaussian spin model, describes the joint distribution of the effective inspiral spin parameter $\chi_{\mathrm{eff}}$ and the effective precession spin parameter $\chi_{\mathrm{p}}$ using a bivariate Gaussian with non-zero correlation [64]. Using posterior samples of the spin model hyperparameters from the population inference [11], we compute the GWB energy spectrum for each sample, modeling spins according to the distributions while keeping all other parameters fixed to those in the fiducial model. To account for spin precession, we employ the waveform model IMRPhenomXP [80].
The influence of spin distributions on the stochastic background energy spectrum is notably smaller than that of mass distribution uncertainties. Therefore, we present only the relative difference with respect to the fiducial model in Fig. 4. At low frequencies, both spin models cause amplitude uncertainties of approximately 3%, while at higher frequencies, the background amplitude generally increases, more prominently for the Gaussian spin model. Around 200 Hz, this enhancement can reach up to 10%. We also note that at low frequencies the spectrum amplitude tends to decrease when spin effects are included. This behavior arises because spins are preferentially aligned [11], which accelerates the inspiral and reduces the energy radiated at lower frequencies [81]. In contrast, aligned-spin binaries experience the orbital hang-up effect [82, 83], allowing tighter orbits just before merger and emit more GW energy at higher frequencies compared to non-spinning systems.
Figure 4: Relative change in the amplitude of the BBH GWB, $\Omega_{\mathrm{BBH}}(f)$, due to the inclusion of spin effects. The curves show deviations from the non-spinning fiducial model at the 90% credible level for two spin distribution models, the Default spin model and the Gaussian spin model.
@@ III.3 Impact of higher-order modes
So far we have computed the background energy spectra using only the dominant (2,2) mode of the waveform. Here, we assess the impact of including higher-order modes by employing the waveform model IMRPhenomXPHM. Since higher-order modes depend differently on the binary’s inclination angle than the (2,2) mode, it is necessary to explicity sample over inclination angles when computing $\Omega_{\mathrm{GW}}(f)$ via Eq. 2, rather than relying on an isotropic average. We first evaluate the fiducial model using IMRPhenomXPHM. As shown in Fig. 5, the inclusion of higher-order modes leads to a modest increase in the background amplitude, particularly at higher frequencies near the peak and fall-off of $\Omega_{\mathrm{GW}}(f)$. At lower frequencies, where most BBHs are in the inspiral phase, the effect is negligible. Next, we keep the PLP mass distribution but select a set of hyperparameters that favor asymmetric masses, specifically we choose one with a low value of $\beta_{q}$. Since higher order modes become more prominent in binaries with unequal masses [84, 85], we expect a stronger impact. Comparing the resulting amplitude increase compared to the same mass distribution evaluated without higher order modes, we find that the increase is greater than in the fiducial model, though still relatively small. Furthermore, we consider a model that retains the fiducial mass distribution but incorporates a spin model favoring large spin magnitudes and tilt angles near $90^{\circ}$. Such a configuration results in a population of strongly precessing binaries, where the GW signal is modulated and subject to mode mixing, thereby increasing the relevance of higher-order modes [86, 87]. Comparing the resulting spectrum to the one computed with only the dominant mode, we find a marginal amplitude increase as compared to the fiducial case. This outcome is expected: while precession modulates the waveform, it has a limited effect on the total radiated GW energy.
@@ III.5 BNS equation of state
The EOS describes the relationship between pressure and density in neutron star matter, thereby determining the mass-radius relation and other macroscopic properties such as tidal deformability. As its precise form is still largely unknown, we investigate how different EOS models influence the energy density spectrum. As in the fiducial model, we assume a uniform distribution of component masses in the range from $1\,M_{\odot}$ to $m_{\mathrm{max,TOV}}$, the maximum mass of a neutron star supported by the selected EOS. To explore a broad region of the EOS parameter space, we consider four representative EOS models: 2H [88], HB [89], APR4 [90] and 2B [88]. These span from stiff EOSs, which predict larger neutron star radii and higher maximum masses, to softer EOSs, which yield more compact stars with lower maximum masses. For example, the APR4 EOS supports a maximum mass of $m_{\mathrm{max,TOV}}=2.2\,M_{\odot}$ and its tidal deformability at the reference mass of $1.4\,M_{\odot}$ is $\Lambda_{1.4}=257$. In Fig. 8, we present the resulting energy density spectra for the different EOSs. Two distinct effects can be observed: first, an increase in amplitude at lower frequencies, which is directly related to the maximum mass supported, and second, a shift of the spectral peak to lower frequencies for stiffer EOSs. This effect is mainly due to the fact that larger neutron stars come into contact earlier during the inspiral, leading to a tapering of the waveform. Additionally, Fig. 8 illustrates the impact of different waveform models on $\Omega_{\mathrm{GW}}(f)$. Compared to the BBH waveform model IMRPhenomXAS, the BNS model exibits tapering in the merger and ringdown phases, even when no tidal deformability is assumed. In contrast, the inspiral-only approximant TaylorF2 [91], which terminates at the last stable circular orbit, deviates from the fiducial model at higher frequencies, by omitting contributions from the merger and ringdown. Finally, we note that an additional EOS-related effect not considered here, is the potential contribution of a post-merger signal, which could impact the background in the kHz regime [92].
@@ III.6 NSBH mass models
The spectral shape of the GWB from NSBH mergers is highly sensitive to the assumed upper limit of the black hole mass when the mass distribution is assumed to be log-uniform, as illustrated in Fig. 9. At low frequencies, we observe the characteristic $f^{2/3}$ scaling of inspiral-dominated signals, with its amplitude governed by the population-averaged chirp mass, as discussed previously. For instance, setting the upper mass limit to $100\,M_{\odot}$, leads to a flattening of the spectrum already at around 10 Hz and it can even peak around 30 Hz due to the increased contribution from high-mass systems merging at lower frequencies. In contrast, when the maximum mass is restricted to $10\,M_{\odot}$, the inspiral regime extends to higher frequencies, resulting in a steeper spectrum and a substantially enhanced amplitude towards 1 kHz. The influence of the neutron star mass distribution on the NSBH energy density spectrum is comparatively minor. An important caveat not addressed here, is the role of tidal interactions, which can significantly affect the GW energy emitted [93, 94]. In particular, tidal disruption of the neutron star may occur when the masses of the black hole and the neutron star are not too unequal, a condition that depends sensitively on the neutron star’s EOS [95, 96, 97]. In such cases, the neutron star can be disrupted before merger, leading to an earlier termination of the waveform and suppression of high-frequency GW emission [98]. As a result, using a BBH waveform model, may overestimate the contribution at higher frequencies to the GWB. Conversely, for large mass ratios, the neutron star typically plunges intact into the black hole, and the waveform is well approximated by a BBH model [99, 100].
@@ IV.1 Time Delay
We start by investigating the effect of the delay time distribution on the resulting merger rate. In Eq. 6 we have introduced the calculation of the merger rate from the SFR convolved with a distribution of time delays $p(t_{d})$. For the fiducial model we have assumed $p\sim t_{d}^{-1}$, which is motivated by the gravitational wave inspiral time scale assuming the initial orbital separation of binaries is flat in log-space [101]. To reflect the uncertainty in the time delay distribution, we examine $p(t_{d})\sim t_{d}^{\nu}$ for $\nu\in[-1.5,-0.5]$. The other critical parameter is the minimum delay time $t_{\mathrm{d,min}}$, which for the fiducial model was assumed to be 50 Myr. In the literature [102, 53, 103, 104, 105, 106, 107], values for $t_{\mathrm{d,min}}\in[10,100]$ Myr are common, depending on the formation channel and assumptions in stellar evolution and population synthesis modeling.
The first column of Fig. 10 displays how extreme choices for the parameters $\nu$ and $t_{\mathrm{d,min}}$ affect the fiducial BBH merger rate (without metallicity cut) and the corresponding background energy density spectrum. A steeper delay-time distribution ($\nu<-1$) favors shorter delay times, leading to a higher peak merger rate and an enhanced background amplitude. Conversely, a flatter distribution or a larger minimal delay time tends to suppress the peak merger rate, which also shifts toward lower redshifts. The relative difference in background amplitude $\Omega_{\mathrm{GW}}$ at 10 Hz between the extreme cases ($\nu=-3/2,\,t_{\mathrm{d,min}}=10\,\mathrm{Myr}$ and $\nu=-1/2,\,t_{\mathrm{d,min}}=100\,\mathrm{Myr}$) exceeds 130%, more than doubling the amplitude. Notably, the shape of the merger rate at low redshifts is particularly sensitive to the delay-time distribution. With upcoming observing runs, we anticipate more precise measurements of the local BBH merger rate and the low-redshift spectral index $\gamma$, which will help constrain the time delay models.
@@ IV.2 Metallicity-weighted SFR
In the past different SFR have been used to compute the binary merger rate and then the resulting background energy density spectrum. Here we explicitly show on three examples that the form of the underlying SFR has a small effect compared to the delay time distribution or metallicity cuts. We compare the Madau-Fragos SFR [67] introduced in Eq. 3 to the Madau-Dickinson SFR [66], which has the same functional form but slightly different parameter values, $\gamma=2.7$, $\kappa=5.6$ and $z_{\mathrm{peak}}=1.9$. Additionally, we consider the SFR provided in Ref. [69] (Vangioni), which is based on the gamma-ray burst rate and is parametrized by the following form given in Ref. [68],
@@
it was used for example in the computation of the CBC background of Ref. [18]. Here we consider the following values: $z_{m}=1.72$, $a=2.8$ and $b=2.46$. In the middle panel of Fig. 10 we illustrate the impact of different SFR models on the BBH merger rate and the resulting GW energy density spectrum, both with and without the metallicity cut defined in Eq. 7. While the Vangioni SFR model predicts a higher merger rate at large redshifts, its influence on the energy density spectrum is minimal. Similarly, the minor changes in the parameter values from the Madau-Dickinson to the Madau-Fragos model do not significantly alter the background.
@@ IV.3 Metallicity model
Here we investigate a particular metallicity model, to compute the merger rate of compact binaries from the metallicity-weighted SFR [103, 66]. We adopt a prescription of the mean metallicity evolution of the Universe based on the assumption that the average metallicity of star-forming gas increases with cosmic time, reflecting the cumulative enrichment by successive generations of stars. The mean metallicity at each redshift is given by
@@
where $R=0.27$ is the return fraction, a net metal yield $y=0.019$, a baryon density $\rho_{b}=2.77\times 10^{11}\Omega_{b}h^{2}\,M_{\odot}\mathrm{Mpc}^{-3}$ with $\Omega_{b}h^{2}=0.0224$, and the specific Madau-Fragos SFR $R_{f}(z)$ [67]. At each redshift we assume a log-normal distribution of metallicity around the mean, with standard deviation $\sigma_{Z}$. Since BBH form preferentially in low-metallicity environments, we can introduce a threshold metallicity $Z_{\mathrm{th}}$ beyond which no BBH formation is possible. The fraction of stars $\varepsilon_{Z}(z)$ formed at redshift $z$ with metallicities below $Z_{\mathrm{th}}$ can be expressed as
@@
In the right panel of Fig. 10, we illustrate how varying the threshold metallicity $Z_{\mathrm{th}}$ and the metallicity dispersion $\sigma_{Z}$ influences the merger rate, computed under the fiducial time delay model. We also show the corresponding impact the GWB energy-density spectrum, comparing results to both the fiducial metallicity model (Equation 7) and the case where no metallicity cut is applied. Introducing a metallicity cut generally shifts the peak of the merger rate to higher redshifts, allowing for larger maximum rates. This shift, in turn, enhances the predicted energy density spectrum of the GWB. Specifically, a low threshold $Z_{\mathrm{th}}$ combined with a narrow dispersion $\sigma_{Z}$ around the mean metallicity leads to elevated merger rates at earlier cosmic times. This occurs because in this case the local Universe contains relatively little low-metallicity star formation, thus, local mergers predominantly originate from the long time-delay tail of a high-redshift progenitor population. As a result, the GWB spectrum exhibits a higher overall normalization and a modified spectral shape near its peak, due to the redshifting of the dominant high-$z$ contribution to lower frequencies.
Figure 10: Impact of different modeling choices on the BBH merger rate and the resulting GWB energy density spectrum. Left panel: Effect of varying the time delay distribution. Shorter minimum delay times and steeper distributions lead to higher merger rates and enhanced background amplitudes. Middle panel: Influence of different SFR models. We compare the merger rates derived from the Madau-Fragos [67], Madau-Dickinson [66] and Vangioni [69] SFR models, shown both with (dashed lines) and without (solid lines) the fiducial metallictiy cut. While the choice of SFR has a minor effect, applying the metallicity cut substantially enhances the merger rate and the resulting GWB amplitude. Right panel: Effect of metallicity modeling. We compare the fiducial metallicity model using two different thresholds with the metallicity model from Ref. [103] evaluated for two thresholds, each with two values of metallicity dispersion. For reference we also show the case without any metallicity cut. Stricter metallicity thresholds and narrower dispersions shift the peak of the merger rate to higher redshifts and increase its amplitude, thereby enhancing the predicted GW background.

### v3

@@ III.2 BBH spin models
We investigate the impact of different spin models on the BBH background energy spectrum and compare them to the nonspinning fiducial models. The GWTC-3 population analysis considered two spin models [11]. The first, referred to as the default model, treats spin magnitudes and orientations independently. Spin magnitudes for both black holes are drawn from a Beta distribution, while tilt angles, defined as the angle between each black hole’s spin and the orbital angular momentum, are drawn from a mixture of an isotropic distribution and a preferentially aligned component, modeled by a truncated Gaussian distribution [79]. The second, known as the Gaussian spin model, describes the joint distribution of the effective inspiral spin parameter $\chi_{\mathrm{eff}}$ and the effective precession spin parameter $\chi_{\mathrm{p}}$ using a bivariate Gaussian with nonzero correlation [64]. Using posterior samples of the spin model hyperparameters from the population inference [11], we compute the GWB energy spectrum for each sample, modeling spins according to the distributions while keeping all other parameters fixed to those in the fiducial model. To account for spin precession, we employ the waveform model IMRPhenomXP [80].
The influence of spin distributions on the stochastic background energy spectrum is notably smaller than that of mass distribution uncertainties. Therefore, we present only the relative difference with respect to the fiducial model in Fig. 4. At low frequencies, both spin models cause amplitude uncertainties of approximately 3%, while at higher frequencies, the background amplitude generally increases, more prominently for the Gaussian spin model. Around 200 Hz, this enhancement can reach up to 10%. We also note that at low frequencies the spectrum amplitude tends to decrease when spin effects are included. This behavior arises because spins are preferentially aligned [11], which accelerates the inspiral and reduces the energy radiated at lower frequencies [81]. In contrast, aligned-spin binaries experience the orbital hang-up effect [82, 83], allowing tighter orbits just before merger and emit more GW energy at higher frequencies compared to nonspinning systems.
Figure 4: Relative change in the amplitude of the BBH GWB, $\Omega_{\mathrm{BBH}}(f)$, due to the inclusion of spin effects. The curves show deviations from the nonspinning fiducial model at the 90% credible level for two spin distribution models, the default spin model and the Gaussian spin model.
@@ III.3 Impact of higher-order modes
So far we have computed the background energy spectra using only the dominant (2,2) mode of the waveform. Here, we assess the impact of including higher-order modes by employing the waveform model IMRPhenomXPHM. Since higher-order modes depend differently on the binary’s inclination angle than the (2,2) mode, it is necessary to explicitly sample over inclination angles when computing $\Omega_{\mathrm{GW}}(f)$ via Eq. 2, rather than relying on an isotropic average. We first evaluate the fiducial model using IMRPhenomXPHM. As shown in Fig. 5, the inclusion of higher-order modes leads to a modest increase in the background amplitude, particularly at higher frequencies near the peak and falloff of $\Omega_{\mathrm{GW}}(f)$. At lower frequencies, where most BBHs are in the inspiral phase, the effect is negligible. Next, we keep the PLP mass distribution but select a set of hyperparameters that favor asymmetric masses, specifically, we choose one with a low value of $\beta_{q}$. Since higher-order modes become more prominent in binaries with unequal masses [84, 85], we expect a stronger impact. Comparing the resulting amplitude increase to the same mass distribution evaluated without higher order-modes, we find that the increase is greater than in the fiducial model, though still relatively small. Furthermore, we consider a model that retains the fiducial mass distribution but incorporates a spin model favoring large-spin magnitudes and tilt angles near $90^{\circ}$. Such a configuration results in a population of strongly precessing binaries, where the GW signal is modulated and subject to mode mixing, thereby increasing the relevance of higher-order modes [86, 87]. Comparing the resulting spectrum to the one computed with only the dominant mode, we find a marginal amplitude increase as compared to the fiducial case. This outcome is expected: while precession modulates the waveform, it has a limited effect on the total radiated GW energy.
@@ III.5 BNS equation of state
The EOS describes the relationship between pressure and density in neutron star matter, thereby determining the mass-radius relation and other macroscopic properties, such as tidal deformability. As its precise form is still largely unknown, we investigate how different EOS models influence the energy density spectrum. As in the fiducial model, we assume a uniform distribution of component masses in the range from $1\,M_{\odot}$ to $m_{\mathrm{max,TOV}}$, the maximum mass of a neutron star supported by the selected EOS. To explore a broad region of the EOS parameter space, we consider four representative EOS models: 2H [88], HB [89], APR4 [90], and 2B [88]. These span from stiff EOSs, which predict larger neutron star radii and higher maximum masses, to softer EOSs, which yield more compact stars with lower maximum masses. For example, the APR4 EOS supports a maximum mass of $m_{\mathrm{max,TOV}}=2.2\,M_{\odot}$ and its tidal deformability at the reference mass of $1.4\,M_{\odot}$ is $\Lambda_{1.4}=257$. In Fig. 8, we present the resulting energy density spectra for the different EOSs. Two distinct effects can be observed: first, an increase in amplitude at lower frequencies, which is directly related to the maximum mass supported, and second, a shift of the spectral peak to lower frequencies for stiffer EOSs. This effect is mainly due to the fact that larger neutron stars come into contact earlier during the inspiral, leading to a tapering of the waveform. Additionally, Fig. 8 illustrates the impact of different waveform models on $\Omega_{\mathrm{GW}}(f)$. Compared to the BBH waveform model IMRPhenomXAS, the BNS model exhibits tapering in the merger and ringdown phases, even when no tidal deformability is assumed. In contrast, the inspiral-only approximant TaylorF2 [91], which terminates at the last stable circular orbit, deviates from the fiducial model at higher frequencies by omitting contributions from the merger and ringdown. Finally, we note that an additional EOS-related effect, not considered here, is the potential contribution of a postmerger signal, which could impact the background in the kHz regime [92].
@@ III.6 NSBH mass models
The spectral shape of the GWB from NSBH mergers is highly sensitive to the assumed upper limit of the black hole mass when the mass distribution is assumed to be log-uniform, as illustrated in Fig. 9. At low frequencies, we observe the characteristic $f^{2/3}$ scaling of inspiral-dominated signals, with its amplitude governed by the population-averaged chirp mass, as discussed previously. For instance, setting the upper mass limit to $100\,M_{\odot}$ leads to a flattening of the spectrum already at around 10 Hz, and it can even peak around 30 Hz due to the increased contribution from high-mass systems merging at lower frequencies. In contrast, when the maximum mass is restricted to $10\,M_{\odot}$, the inspiral regime extends to higher frequencies, resulting in a steeper spectrum and a substantially enhanced amplitude toward 1 kHz. The influence of the neutron star mass distribution on the NSBH energy density spectrum is comparatively minor. An important caveat not addressed here is the role of tidal interactions, which can significantly affect the GW energy emitted [93, 94]. In particular, tidal disruption of the neutron star may occur when the masses of the black hole and the neutron star are not too unequal, a condition that depends sensitively on the neutron star’s EOS [95, 96, 97]. In such cases, the neutron star can be disrupted before merger, leading to an earlier termination of the waveform and suppression of high-frequency GW emission [98]. As a result, using a BBH waveform model may overestimate the contribution at higher frequencies to the GWB. Conversely, for large mass ratios, the neutron star typically plunges intact into the black hole, and the waveform is well approximated by a BBH model [99, 100].
@@ IV.1 Time delay
We start by investigating the effect of the time-delay distribution on the resulting merger rate. In Eq. 6 we have introduced the calculation of the merger rate from the SFR convolved with a distribution of time delays $p(t_{d})$. For the fiducial model we have assumed $p\sim t_{d}^{-1}$, which is motivated by the gravitational wave inspiral timescale assuming the initial orbital separation of binaries is flat in log-space [101]. To reflect the uncertainty in the time-delay distribution, we examine $p(t_{d})\sim t_{d}^{\nu}$ for $\nu\in[-1.5,-0.5]$. The other critical parameter is the minimum delay time $t_{\mathrm{d,min}}$, which for the fiducial model was assumed to be 50 Myr. In the literature [102, 53, 103, 104, 105, 106, 107], values for $t_{\mathrm{d,min}}\in[10,100]$ Myr are common, depending on the formation channel and assumptions in stellar evolution and population synthesis modeling.
The first column of Fig. 10 displays how extreme choices for the parameters $\nu$ and $t_{\mathrm{d,min}}$ affect the fiducial BBH merger rate (without metallicity cut) and the corresponding background energy density spectrum. A steeper delay-time distribution ($\nu<-1$) favors shorter delay times, leading to a higher peak merger rate and an enhanced background amplitude. Conversely, a flatter distribution or a larger minimal delay time tends to suppress the peak merger rate, which also shifts toward lower redshifts. The relative difference in background amplitude $\Omega_{\mathrm{GW}}$ at 10 Hz between the extreme cases ($\nu=-3/2,\,t_{\mathrm{d,min}}=10\,\mathrm{Myr}$ and $\nu=-1/2,\,t_{\mathrm{d,min}}=100\,\mathrm{Myr}$) exceeds 130%, more than doubling the amplitude. Notably, the shape of the merger rate at low redshifts is particularly sensitive to the delay-time distribution. With upcoming observing runs, we anticipate more precise measurements of the local BBH merger rate and the low-redshift spectral index $\gamma$, which will help constrain the time-delay models.
@@ IV.2 Metallicity-weighted SFR
In the past, different SFRs have been used to compute the binary merger rate and then the resulting background energy density spectrum. Here, we explicitly show in three examples that the form of the underlying SFR has a small effect compared to the delay-time distribution or metallicity cuts. We compare the Madau-Fragos SFR [67] introduced in Eq. 3 to the Madau-Dickinson SFR [66], which has the same functional form but slightly different parameter values, $\gamma=2.7$, $\kappa=5.6$, and $z_{\mathrm{peak}}=1.9$. Additionally, we consider the SFR provided in Ref. [69] (Vangioni), which is based on the gamma-ray burst rate and is parametrized by the following form given in Ref. [68]:
@@
which was used, for example, in the computation of the CBC background of Ref. [18]. Here, we consider the following values: $z_{m}=1.72$, $a=2.8$, and $b=2.46$. In the middle panel of Fig. 10 we illustrate the impact of different SFR models on the BBH merger rate and the resulting GW energy density spectrum, both with and without the metallicity cut defined in Eq. 7. While the Vangioni SFR model predicts a higher merger rate at large redshifts, its influence on the energy density spectrum is minimal. Similarly, the minor changes in the parameter values from the Madau-Dickinson to the Madau-Fragos model do not significantly alter the background.
@@ IV.3 Metallicity model
Here, we investigate a particular metallicity model to compute the merger rate of compact binaries from the metallicity-weighted SFR [103, 66]. We adopt a prescription of the mean metallicity evolution of the Universe based on the assumption that the average metallicity of star-forming gas increases with cosmic time, reflecting the cumulative enrichment by successive generations of stars. The mean metallicity at each redshift is given by
@@
where $R=0.27$ is the return fraction, a net metal yield $y=0.019$, a baryon density $\rho_{b}=2.77\times 10^{11}\Omega_{b}h^{2}\,M_{\odot}\mathrm{Mpc}^{-3}$ with $\Omega_{b}h^{2}=0.0224$, and the specific Madau-Fragos SFR $R_{f}(z)$ [67]. At each redshift we assume a log-normal distribution of metallicity around the mean, with standard deviation $\sigma_{Z}$. Since BBHs form preferentially in low-metallicity environments, we can introduce a threshold metallicity $Z_{\mathrm{th}}$ beyond which no BBH formation is possible. The fraction of stars $\varepsilon_{Z}(z)$ formed at redshift $z$ with metallicities below $Z_{\mathrm{th}}$ can be expressed as
@@
In the right panel of Fig. 10, we illustrate how varying the threshold metallicity $Z_{\mathrm{th}}$ and the metallicity dispersion $\sigma_{Z}$ influences the merger rate, computed under the fiducial time-delay model. We also show the corresponding impact on the GWB energy-density spectrum, comparing results to both the fiducial metallicity model [Eq. 7] and the case where no metallicity cut is applied. Introducing a metallicity cut generally shifts the peak of the merger rate to higher redshifts, allowing for larger maximum rates. This shift, in turn, enhances the predicted energy density spectrum of the GWB. Specifically, a low threshold $Z_{\mathrm{th}}$ combined with a narrow dispersion $\sigma_{Z}$ around the mean metallicity leads to elevated merger rates at earlier cosmic times. This occurs because in this case, the local Universe contains relatively little low-metallicity star formation, thus, local mergers predominantly originate from the long time-delay tail of a high-redshift progenitor population. As a result, the GWB spectrum exhibits a higher overall normalization and a modified spectral shape near its peak, due to the redshifting of the dominant high-$z$ contribution to lower frequencies.
Figure 10: Impact of different modeling choices on the BBH merger rate and the resulting GWB energy density spectrum. Left panel: effect of varying the time-delay distribution. Shorter minimum delay times and steeper distributions lead to higher merger rates and enhanced background amplitudes. Middle panel: influence of different SFR models. We compare the merger rates derived from the Madau-Fragos [67], Madau-Dickinson [66], and Vangioni [69] SFR models, shown both with (dashed lines) and without (solid lines) the fiducial metallictiy cut. While the choice of SFR has a minor effect, applying the metallicity cut substantially enhances the merger rate and the resulting GWB amplitude. Right panel: effect of metallicity modeling. We compare the fiducial metallicity model using two different thresholds with the metallicity model from Ref. [103] evaluated for two thresholds, each with two values of metallicity dispersion. For reference we also show the case without any metallicity cut. Stricter metallicity thresholds and narrower dispersions shift the peak of the merger rate to higher redshifts and increase its amplitude, thereby enhancing the predicted GW background.